### v1

# Minds versus Machines: Rethinking Entailment Verification with Language Models

Abstract
Humans make numerous inferences in text comprehension to understand discourse. This paper aims to understand the commonalities and disparities in the inference judgments between humans and state-of-the-art Large Language Models (LLMs). Leveraging a comprehensively curated entailment verification benchmark, we evaluate both human and LLM performance across various reasoning categories. Our benchmark includes datasets from three categories (NLI, contextual QA, and rationales) that include multi-sentence premises and different knowledge types, thereby evaluating the inference capabilities in complex reasoning instances. Notably, our findings reveal LLMs’ superiority in multi-hop reasoning across extended contexts, while humans excel in tasks necessitating simple deductive reasoning. Leveraging these insights, we introduce a fine-tuned Flan-T5 model [^1] https://huggingface.co/soumyasanyal/entailment-verifier-xxl that outperforms GPT-3.5 and rivals with GPT-4, offering a robust open-source solution for entailment verification. As a practical application, we showcase the efficacy of our finetuned model in enhancing self-consistency in model-generated explanations, resulting in a 6% performance boost on average across three multiple-choice question-answering datasets.
Minds versus Machines: Rethinking Entailment Verification with Language Models
Soumya Sanyal1111 Tianyi Xiao1111 Jiacheng Liu2222 Wenya Wang3333Xiang Ren1111 $1$University of Southern California $2$University of Washington $3$Nanyang Technological University, Singapore soumyasa@usc.edu
[…]
[sections/figures.tex]motivation
[…]
[sections/tables.tex]datasetComparisons
1 Introduction
A prevailing notion in cognitive psychology exists that humans make numerous inferences to understand discourse and text Garnham (1989). These inferences, serving diverse functions in text comprehension, extend beyond resolving lexical ambiguities or determining the referent of a pronoun; they play a crucial role in linking information from disparate sections of a text to establish its literal meaning. Cognitive studies Buschman et al. (2011); Cowan (2001) have shown that an average human brain has a limited capacity to retain only four chunks in short-term memory, indicating a limitation of human inference abilities. Similarly, prior works Gururangan et al. (2018); McCoy et al. (2019); Geiger et al. (2020); Clark et al. (2020); Sanyal et al. (2022) have shown that large language models (LLMs) are still subpar at understanding negations in inference, use spurious correlations while reasoning, cannot reason through multi-step compositions robustly, etc. With the advent of large language models (LLMs) and their applications on natural language inference (NLI) Minaee et al. (2021); Nie et al. (2020), understanding the key difference in the inference abilities between humans and LLMs is crucial to making further progress, which is currently missing.
In this work, we aim to address this by first evaluating both humans and current LLMs Liu et al. (2019); Tafjord et al. (2022); Chung et al. (2022); Brown et al. (2020); OpenAI (2023) on a comprehensive entailment verification benchmark and then analyzing the performance across different reasoning categories based on the ease of inference deduction and the type of knowledge required. Existing textual inference datasets such as SNLI Bowman et al. (2015), MNLI Williams et al. (2018), etc., mostly contain short sentence premises that do not completely encapsulate the challenges of multi-sentence, complex reasoning. Predicting the entailment of such complex premise-hypothesis pairs often requires multi-hop reasoning, inferring missing information, robustness to spurious correlation, etc., which are largely missing from simpler datasets Gururangan et al. (2018); McCoy et al. (2019). Also, real-world scenarios that require inference skills, such as understanding stories, dialogues, etc., usually contain multi-sentence premises. Therefore, we select multiple datasets across three categories (NLI, contextual QA, and rationales) and convert them into NLI format, as required, to create the evaluation benchmark. As shown in Table LABEL:tab:dataset_comparisons, the datasets used in our study typically contain multi-sentence premises that require different types of knowledge to predict the entailment.
Our analyses show that LLMs are stronger than humans at entailment tasks that involve multi-hop reasoning across long contexts. This suggests that training LLMs to reason on diverse long context data is beneficial for entailment verification on complex cases. In contrast, we find that humans are better at solving entailment tasks that require simple deductive reasoning using substitutions, negations, etc., indicating that current LLMs still lack consistency along these reasoning aspects. Further, we find that humans and LLMs perform comparably in entailment tasks requiring inferring some missing knowledge. These findings are further depicted in Figure LABEL:fig:motivation with motivating examples.
Among the different LLMs, we find that general instruction-finetuned models are better than task-finetuned models trained on a specific dataset category. We leverage this finding by finetuning a Flan-T5 Chung et al. (2022) model on a training subset containing datasets from each category. We explore two different training approaches: a classification-based finetuning that learns to directly predict the label and a ranking-based finetuning that learns to rank the most supported hypothesis from a given pair of hypotheses for a given premise. We find that ranking-based finetuning is superior to classification as it can learn a softer decision boundary. Overall, our fine-tuned models outperform GPT-3.5 and perform comparably to GPT-4 on the benchmark, thus providing a strong open-sourced model for entailment verification.
Finally, we demonstrate the utility of our finetuned models on a downstream application of filtering unfaithful model-generated explanations. Recent works Wei et al. (2022); Zhou et al. (2023); Wang et al. (2023) have prompted LLMs to generate natural language explanations for model predictions, but such reasoning steps can sometimes be inconsistent, i.e., the explanation does not entail the model’s final prediction Ye and Durrett (2022). In self-consistency (SC) Wang et al. (2023), multiple reasoning paths are first sampled from the LLM decoder for a given instance and then aggregated to predict the most consistent answer. We use our finetuned models to filter out non-entailed reasoning chains in the SC decoding strategy before aggregating the final prediction, which leads to 6% performance improvement on average across three MCQ datasets.
2 Humans vs. LLMs on Entailment Verification
In this section, we define the task of entailment verification (EV), datasets used to create the evaluation benchmark, the evaluation procedure for evaluating different LLM baselines and humans, and compare them across different reasoning types.
[…]
2.1 Evaluation Benchmark
To benchmark humans and LLMs on the task of entailment verification, a basic criterion for dataset selection is that the dataset should contain a premise, a hypothesis, and a support/not support label associated with it. Additionally, we list some more desirable characteristics we want to cover in this study, as discussed below:
  • Type of Premise: Typically, NLI datasets, such as SNLI, MNLI, etc., do not contain more than one sentence in the premise, potentially leading to shortcut learning. In contrast, we focus more on multi-sentence premises that require complex reasoning. We also consider datasets where the premise is a rationale, i.e., the premise is not just a logical precursor to the hypothesis but rather an explanation. This tests the ability to evaluate model-generated rationales Wei et al. (2022).
  • Type of Knowledge: Often, one or more information in the premise needs to be used to predict support. We categorize this information as entity-grounded, commonsense, or localized. Entity-grounded knowledge consists of information about entities and other general knowledge verifiable on the internet. These can be facts about general science, history, etc., or details of some known person, event, etc. It is possible to infer this information even if not mentioned in the premise. The commonsense knowledge is typically all information about everyday life that humans use implicitly but cannot always be verified online. This information is often missing from the premise and has to be inferred implicitly. Lastly, localized information is all other knowledge provided for understanding the events, people, or items mentioned in the premise that is not grounded to any known entity. This information depends on the premise’s specific context and, thus, is impossible to infer unless stated explicitly. Please refer to Table LABEL:tab:knowledge_examples for examples of each knowledge type.
[…]
[sections/tables.tex]knowledgeExamples
We consider three data sources for creating the entailment verification benchmark, amounting to 10 datasets in total. In Table LABEL:tab:dataset_comparisons, we compare these datasets across the desirable characteristics mentioned earlier. Please refer to Appendix A for more details on the datasets used.
[…]
[sections/tables.tex]humanResults
Natural Language Inference
NLI is an obvious choice of data source. While converting an NLI dataset for our task, we merge the neutral and contradict labels to the not support label. We use the following NLI datasets in our benchmark: WaNLI Liu et al. (2022), FEVER Nie et al. (2019), and ANLI Nie et al. (2020).
Contextual QA
Next, we consider QA datasets where the task is to answer a question based on a given context and some options. We use an off-the-shelf QA-to-statement converter model Chen et al. (2021) to generate a hypothesis statement for each question option pair. Then, the hypothesis corresponding to the correct choice is marked as “support”, while the rest are marked as “not support” to create the entailment verification dataset. Overall, we include the following datasets from this category: Cosmos QA (CosQA) Huang et al. (2019), SocialIQA (SIQA) Sap et al. (2019), DREAM Sun et al. (2019), BoolQ Clark et al. (2019), and RACE Lai et al. (2017).
Rationale
Lastly, we consider data sources where human-annotated explanations are available that justify the original hypothesis (or the correct option, in the case of QA datasets). In this case, we use the rationales as the premise. We use the following datasets: Entailer Tafjord et al. (2022), and ECQA Aggarwal et al. (2021).
2.2 LLM Evaluation Setup
We evaluate two types of LLMs on the task of entailment verification, as categorized below:
Task-finetuned LLMs
In this category, the models considered are already finetuned for either NLI or the exact entailment verification task itself. We evaluate RoBERTa (finetuned on NLI data) and Entailer-11B (finetuned on Entailer data) in this category. Please refer to Appendix B.1 for more details on the evaluation setup for these models.
Instruction-finetuned LLMs
These are “general-purpose” language models trained on a collection of NLP tasks described using instructions, leading to generalization abilities to solve unseen tasks described using new instructions. Here, we evaluate Flan-T5-xxl, GPT-3.5, and GPT-4 models. To compute the label, we first modify a given premise-hypothesis pair $(p,h)$ into a prompted input $\mathcal{P}$ using the prompt template as shown in Box 1. Next, we compute a score $s$ as defined below:
$s(p,h)=\frac{p_{LLM}({\textrm{``Yes''}|\mathcal{P}})}{p_{LLM}({\textrm{``Yes''}|\mathcal{P}})+p_{LLM}({\textrm{``No''}|\mathcal{P}})},$ (1)
where $p_{LLM}(\cdot|\mathcal{P})$ is the model’s probability distribution over the vocabulary. If the score $s$ is higher than a threshold (typically set to $0.5$ in all our experiments), we assign the label support, else we assign the label not support. For GPT-4 evaluation, we directly check for the “Yes” / “No” label prediction as the token probabilities are not accessible via the API. Please refer to Appendix B.2 for more details about the models and ablations on few-shot prompts.
Premise: {premise} Hypothesis: {hypothesis} Question: Given the premise, is the hypothesis correct? Answer:
List of Boxes 1 Prompt used to evaluate instruction-finetuned LLMs for entailment verification.
2.3 Evaluation Metric
We use the macro-F1 score as the primary evaluation metric for comparing LLMs on the entailment verification task because there are label imbalances in our evaluation datasets. The macro-F1 score computes the unweighted mean of F1 scores for each class, ensuring equal importance for each class irrespective of the label statistics. Please refer to Appendix C for more discussions on the label imbalance of each dataset.
2.4 Human vs. LLMs
First, we randomly sample 100 instances for each dataset (i.e., 1000 instances in total) and conduct a human evaluation on this subset to estimate average human performance. Please refer to Appendix E.1 for more details on the annotation procedure. Additionally, we evaluate the above LLMs on this sampled subset and report those numbers for fair comparisons with humans. Table LABEL:tab:human_results shows the overall evaluation results. Among LLMs, we observe that instruction-finetuned LLMs are stronger than task-finetuned models that are trained on some selected datasets. This shows that using a specific dataset for fine-tuning does not necessarily improve the general inference abilities of LLMs. On comparisons between humans and LLMs, we find that humans beat all the baseline LLMs, except GPT-4. This shows that existing open-sourced LLMs are not at par with humans on this task. Additionally, although humans and GPT-4 perform comparably on average, there are large misalignments in different individual datasets. Specifically, we observe that ANLI, CosQA, SIQA, and Entailer are the four datasets with $>0.1$ absolute macro-F1 difference. Next, we perform more analysis on these four datasets to better understand these misalignments.
[…]
[sections/figures.tex]reasoningAnalysis2
2.5 Reasoning Type Analysis
We design a new analysis to further understand the reasons behind such misalignment between models and humans. First, we categorize the type of reasoning required to successfully predict an entailment into the following four categories:
  • Simple Deductive (R1): In this type, the premise contains sentences that can be minimally combined in one step to predict the support for the hypothesis. This type typically tests skills such as substitution, negation, synonyms, etc.
  • Complex Deductive (R2): In this, more than one step of reasoning using information from the premise is required to solve the task. Typically, this type tests skills like mathematical reasoning, using multiple information in context to arrive at a conclusion, etc.
  • Missing Entity-grounded/Commonsense Knowledge (R3): This is a case where some essential commonsense or entity-grounded knowledge is missing in the premise. This gap can potentially be bridged by the model’s parametric knowledge obtained from pretraining.
  • Missing Localized Knowledge (R4): In this, information very specific to the context of the premise is missing. Typically, this is information about the subjects in the context and is not grounded to any entities that can be known via the internet. It is practically impossible for the model (or even humans) to infer such missing information.
Please refer to Appendix E.3 for examples of each reasoning type. Given this categorization, we first annotate the reasoning type of the 100 previously sampled instances for each dataset with absolute macro-F1 difference $>0.1$. We make this choice intending to attribute the largely misaligned datasets (namely, ANLI, CosQA, SIQA, and Entailer) since some random noise in the annotation and sampling process can potentially also cause some misalignment. Please refer to Appendix E.2 for more details on the annotation setup. Figure LABEL:fig:reasoning_analysis2 depicts the aggregated results for each reasoning type. The top plot shows the frequency of each reasoning type, and the bottom plot compares the human and GPT-4 macro-F1 scores. Please refer to Appendix E.4 for a detailed analysis of individual datasets.
The first type in Figure LABEL:fig:reasoning_analysis2 is simple deductive reasoning ($\sim$ 25% data). Here, humans perform better than GPT-4 by a small margin. Instances that require simple deductive reasoning usually use substitutions, negations, paraphrasing, etc., to prove entailment (refer to Table LABEL:tab:reasoning_examples for examples). We find that humans are more robust than GPT-4 in performing such simple deductive reasoning tasks, as observed in prior works Sanyal et al. (2022); Nguyen et al. (2023).
[…]
[sections/tables.tex]finetuneResults
Next, we find that GPT-4 significantly outperforms the human baseline on complex reasoning that constitutes $\sim$ 20% data. This type of reasoning usually requires two skills: understanding multiple relevant information in the premise and combining them for reasoning. There are two potential reasons for this misalignment. First, compared to humans, GPT-4 is likely a stronger context processor, especially for long premises, since it has been trained on long-context data sources OpenAI (2023). Alternatively, another potential confounding factor is that AMT workers are typically incentivized to quickly finish annotations to maximize earnings. Thus, instances that require multi-step reasoning can have noisier annotations, potentially underestimating true human performance.
Lastly, we observe that approximately 30% of the data has some missing entity-grounded or commonsense information while $\sim$ 25% of the data has missing localized information. To correctly predict entailment in such instances, a system should be able to infer some of the missing grounded knowledge while not hallucinating specific localized information not mentioned in the premise. We find that both humans and models are comparable across the reasoning types R3 and R4, with R4 being more challenging. This shows that both models and humans tend to hallucinate missing localized information.
3 Training LLMs for Entailment Verification
In Section 2.4, we observed that open-sourced LLMs are lacking in performance compared to humans and close-sourced models such as GPT-4. Thus, in this section, we propose training strategies to finetune an open-sourced model that can perform competitively w.r.t. GPT-4. To this end, we finetune Flan-T5-xxl model using the training data from some of the above-mentioned datasets. Specifically, we choose the training portions of ANLI, RACE, and ECQA for finetuning the model. Please refer to Appendix F.1 for more details on our training dataset selection criteria. We describe two approaches to finetuning and then discuss our findings.
[…]
[sections/figures.tex]cotFlowchart
3.1 Finetuning Formulations
This section describes the two fine-tuning formulations explored in this work.
Classification
This is the standard training paradigm where we finetune a Flan-T5-xxl model using the training data. We follow the same steps as the evaluation setup to create a prompted input using the prompt format in Box 1 and then define the cross-entropy loss over the “Yes” and “No” token logits. We refer to this finetuned model as “Flan-T5-xxl + Class”.
Ranking
In this approach, for a given premise and hypothesis pair $(p,h)$, we define a weaker hypothesis $h^{\prime}$ as a statement such that the premise $p$ supports $h$ more strongly than $h^{\prime}$. Then, for a given triple $(p,h,h^{\prime})$, we formulate the ranking task as predicting the hypothesis that is more supported by the premise. Given the triplet $(p,h,h^{\prime})$, we define the margin ranking loss as follows:
$\mathcal{L}_{ranking}=\max\{0,s(p,h)-s(p,h^{\prime})+m\},$ (2)
where $s(p,h)$ is the entailment score as defined in Equation 1. The key advantage of this formulation over the classification is that ranking, by design, learns a softer decision boundary between the two labels. This can lead to better generalization, especially for contextual QA datasets. Sometimes, the wrong choice can be relatively less favorable w.r.t. to the best option in QA instead of being absolutely incorrect. Training to hard-classify the hypothesis for such options can be avoided by ranking them with the best hypothesis (corresponding to the right choice), thus learning a softer classification boundary. We refer to the finetuned model using the ranking objective as “Flan-T5-xxl + Rank”. Please refer to Appendix F.2 for more details on the training data collection process for ranking.
3.2 Findings
Table LABEL:tab:finetune_results shows the evaluation results on the complete evaluation set (i.e., we use all the data points instead of 100 samples per dataset, which was used in Table LABEL:tab:human_results). For our models, we separately average the results for the datasets already seen in training (namely, ANLI, RACE, and ECQA) and unseen during training into two columns, seen and unseen, respectively. First, we observe that all our finetuned models are consistently better across nine out of ten datasets than the baseline Flan-T5-xxl. Finetuning improves 0.07 macro-F1 on average. This shows that finetuning is overall beneficial in training the model on the task of entailment verification.
Next, we observe that compared to classification, the ranking formulation is beneficial for the contextual QA datasets CosQA, SIQA, and DREAM. This demonstrates that the ranking objective improves contextual QA datasets’ generalization, which is expected. Additionally, our ranking model outperforms GPT-3.5 and performs comparably to GPT-4, with stronger performance on contextual QA datasets and weaker performance on NLI datasets. Thus, Flan-T5-xxl + Rank is a strong open-sourced model for entailment verification and can be used as an alternative to GPT-4.
4 Application: Filtering CoT Rationales
Recently, Wang et al. (2023) proposed self-consistency (SC), a decoding technique to improve over chain-of-though (CoT) reasoning Wei et al. (2022) in LLMs, whereby multiple CoT rationales are sampled for a given input instance and a majority voting overall predicted labels is considered as the final prediction. But generative LLMs can potentially output rationales that are inconsistent Ye and Durrett (2022), i.e., the rationale does not support the corresponding model prediction. Such inconsistency can, in turn, degrade the overall self-consistency results. This motivates an important application of entailment verification systems in this paradigm. Please refer to Appendix G.1 for examples of consistent and inconsistent CoTs.
Approach
As shown in Figure LABEL:fig:cot_application, we can use a verifier as an intermediate step before computing the majority vote, whereby the inconsistent rationales are filtered out. We achieve this by first scoring all the rationales (premises) and their corresponding predictions (hypothesis [^2] For MCQ datasets, typically, the prediction is one of the choices. We use Chen et al. (2021) as defined in Section 2.1 to convert the question-answer pair into a statement.) using Equation 1. Then, we use the top-$k$ rationales for majority voting. We set $k=10$ for all our experiments.
[…]
[sections/figures.tex]cot
Findings
In figure LABEL:fig:cot_comparison, we compare the vanilla SC with the filtering+SC approach described above. Following Wang et al. (2023), we compute the average performance of these methods across three MCQ datasets for four different base CoT models (UL2 Tay et al. (2023), Codex-001 Brown et al. (2020), LaMDA-137B Thoppilan et al. (2022), and ChatGPT OpenAI (2022)). Please refer to Appendix G for details on the datasets and more comparisons with Flan-T5-xxl. We observe that filtering leads to a consistent performance gain over SC across all CoT base models. This demonstrates the advantage of the filtering approach. Next, we find that the improvements are more prominent for weaker base models such as UL2 than the stronger ones (ChatGPT). For instance, filtering UL2 generated-rationales can even achieve comparable performance with vanilla SC over LaMDA-137B. In comparison, the gains for filtering ChatGPT CoTs are $\sim$ 1.5 %. This shows that weaker models are prone to generating inconsistent CoTs and thus benefit more from this approach. But at the same time, even stronger models such as ChatGPT can still benefit from consistency checks. Please refer to Appendix G.1 for examples of filtered CoT rationales.
5 Related Works
Natural Language Inference
NLI Dagan et al. (2006); Manning and MacCartney (2009) is one of the core NLP problems in which the relationship between a premise and hypothesis is classified as either entailment, contradiction, or neutral. Prior works have mainly trained LLMs and evaluated them on standard NLI datasets Bowman et al. (2015); Williams et al. (2018); Wang et al. (2019); Nie et al. (2020, 2019); Liu et al. (2022). Another line of work Mishra et al. (2020); Chen et al. (2021) has used question-answer-to-NLI conversion Demszky et al. (2018) to transform QA datasets into NLI format and solve them. In fact verification literature Bekoulis et al. (2021); Thorne et al. (2018), retrieved pieces of evidence have been used to verify the claim using an entailment verifier Nie et al. (2019); Guan et al. (2023). Recently, NLI models have been used to verify the entailment of model-generated explanations Tafjord et al. (2022); Jung et al. (2022); Mitchell et al. (2022). In this work, we curate a diverse NLI benchmark for evaluating LLMs and humans by using datasets from all the above NLI applications.
Reasoning in LLMs
With the advent of strong, general-purpose LLMs Brown et al. (2020); Chung et al. (2022); OpenAI (2023), many prompting strategies have been proposed to generate a natural language reasoning along with the model’s prediction Wei et al. (2022); Zhou et al. (2023); Yao et al. (2023); Huang and Chang (2023). Recently, Ye and Durrett (2022) have found that such generations can sometimes be unreliable due to non-factual and inconsistent reasoning, while Huang et al. (2023) have argued that LLMs struggle to self-correct such issues without external feedback.
Prior works have addressed this limitation by oversampling reasoning chains and marginalizing Wang et al. (2023), using the LLMs itself to recheck their reasoning Madaan et al. (2023); Miao et al. (2023), leveraging external knowledge source to verify factuality Zhao et al. (2023), using deterministic solvers to improve faithfulness Lyu et al. (2023), decomposing the reasoning steps into smaller steps Ling et al. (2023), etc. While the progress is impressive, some of these are either specialized approaches for math-specific datasets or heavily rely on close-sourced LLMs (GPT-3.5, GPT-4, etc.) for verification. In contrast, here we focus on natural language datasets and develop a strong open-sourced LLM that can be easily deployed to check entailment errors in LLM reasoning chains.
6 Conclusion
We studied the entailment verification problem in the context of LLMs. Specifically, we sourced datasets across three different categories (NLI, contextual QA, and rationales) and analyzed both human and model performance across these datasets. We found some misalignments between the two, whereby models are better than humans in complex reasoning and humans are relatively more consistent on simpler reasoning tasks. We also explored different finetuning objectives to train LLMs for verification. Our fine-tuned models outperform GPT-3.5 and are at par with GPT-4. Finally, we demonstrated a practical application of entailment verification in filtering out inconsistent model-generated rationales while using self-consistency decoding. Overall, we believe our work provides a meaningful comparison of inference capabilities between humans and LLMs, a comprehensive evaluation benchmark, an open-sourced entailment verification model, and interesting applications of the task.
Limitations
Even though our work demonstrates exciting results on entailment verification tasks by finetuning LLMs, several limitations can be potentially improved. We only tried encoder-decoder-based models for finetuning. However, other models with different architectures (like decoder-only) can also be considered. Our strategy to convert a QA pair into a statement using the QA-to-statement converter model can have errors that can cascade both in the evaluation dataset and our fine-tuned models. For computing the entailment score in Equation 1, we only considered the probability of “Yes” and “No” tokens in the entire vocabulary. Other alternative expressions like “YES”/“NO”, “True”/“False”, etc., can also be considered to make the score more robust. Finally, our training objective outputs entailment scores instead of directly generating answers. Answer generation as a training objective can be more robust since it is a stricter objective than our scoring technique.
Acknowledgments
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the DARPA MCS program under Contract No. N660011924033, the Defense Advanced Research Projects Agency with award W911NF-19-20271, NSF IIS 2048211, and gift awards from Google and Amazon. The views and conclusions contained herein are those of the authors. They should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. We would like to thank all the USC INK research lab collaborators for their constructive feedback on the work.
Appendix A Evaluation Datasets
In this section, we describe the datasets used in our evaluation. We mention some important challenges that make these datasets useful for benchmarking the entailment verification task. Please refer to Table LABEL:tab:dataset_statistics for these datasets’ train/dev/test statistics.
[…]
[sections/tables.tex]dataStatistic
Natural Language Inference datasets
NLI is an obvious choice of data source as it is a more general case of the entailment verification problem. While converting an NLI dataset for our task, we merge the neutral and contradict labels to the not support label. We use the following NLI datasets in our benchmark:
  • WaNLI Liu et al. (2022): This is a new NLI dataset built using worker and AI collaboration. This challenging dataset improves over the existing NLI dataset MultiNLI Williams et al. (2018). And the test split is used when doing evaluation.
  • FEVER Nie et al. (2019): This is a modification of the original FEVER dataset Thorne et al. (2018) in which the claim is paired with textual evidence from Wikipedia to convert it into an NLI format dataset. This pairing uses existing state-of-the-art evidence extraction systems to find relevant evidence for each claim. Premises in this dataset typically contain multiple sentences, which is one of our focus areas. As the test split is not available, we report results on dev split for evaluation.
  • ANLI Nie et al. (2020): This is a large-scale NLI dataset that was collected using an adversarial human-and-model-in-the-loop procedure. Like FEVER, this dataset tests factual knowledge, and the premises typically contain multiple sentences. During evaluation, the test split is considered.
Contextual QA datasets
Next, we consider QA datasets where the task is to answer a question based on a given context and some options. We use an off-the-shelf QA-to-statement converter model Chen et al. (2021) to generate a hypothesis statement for each question option pair. Then, the hypothesis corresponding to the correct choice is marked as “support”, while the rest are marked as “not support” to create the entailment verification dataset. Overall, we include the following datasets from this category:
  • Cosmos QA (CosQA) Huang et al. (2019): This dataset contains multiple-choice questions (MCQs) that require an understanding of commonsense-based reading comprehension to answer a question. The key challenge in this dataset is understanding people’s everyday narratives described in the context that can have some missing commonsense knowledge that needs to be inferred implicitly. Since the test split is missing, we evaluate models on dev split instead.
  • SocialIQA (SIQA) Sap et al. (2019): Similar to CosQA, this is another MCQ benchmark for commonsense reasoning about social situations that probes emotional and social intelligence in a variety of everyday situations. This dataset has more nuanced commonsense knowledge requirements, which makes it a challenging dataset for our task. Similarly, results on dev split are reported given test is missing.
  • DREAM Sun et al. (2019): This is a dialogue-based reading comprehension MCQ dataset that focuses on multi-turn dialogue understanding. Here, the unique challenge is inferring the events discussed across long, multi-turn dialogues. During evaluation, we use the test split as it is available.
  • BoolQ Clark et al. (2019): This is a True/False QA dataset consisting of aggregated queries to the Google search engine. Questions in this dataset require complex and difficult entailment-like inference to solve, making it a good set for evaluation. The test split is lacking for this dataset and we can only report results on dev split.
  • RACE Lai et al. (2017): The reading comprehension dataset from examinations (RACE) is one of the most popular machine reading comprehension datasets containing questions from English exams for middle and high school students. These questions are designed by domain experts for testing specific human reading skills, thus making it a good evaluation set for our task. We report evaluating results on test split for this dataset.
Rationale datasets
Lastly, we consider data sources where human-annotated explanations are available that justify the original hypothesis (or the correct option, in the case of QA datasets). In this case, we use the rationales as the premise. We use the following datasets:
  • Entailer Tafjord et al. (2022): This dataset contains entailment-style statements and corresponding rationales obtained from EntailmentBank dataset Dalvi et al. (2021) and crowdsourcing. The dataset mainly contains science domain statements and tests simple deductive reasoning skills whereby sentences from the premise have to be combined to either support or refute the hypothesis. The test split is also missing for this dataset and we can only evaluate on the dev split.
  • ECQA Aggarwal et al. (2021): This is a human-annotated explanation dataset for CommonsenseQA Talmor et al. (2019). We only use the explanations for the correct choice as the explanations for the incorrect choices are often trivial. It is a complete dataset and, by convention, we use the test split for evaluation.
Appendix B Details on LLM Evaluation
We evaluate two types of LLMs on the task of entailment verification, as categorized below:
B.1 Task-Finetuned models
In this category, the models considered are already finetuned for either NLI or the exact entailment verification task itself. We evaluate two models in this category.
RoBERTa
Nie et al. (2020); Liu et al. (2019) This is a strong pre-trained RoBERTa-Large model with corresponding model card on HuggingFace [^3] https://huggingface.co/models called “ynie/roberta-large-snli_mnli_fever_anli_R1_R2_R3-nli”. It is a specifically pre-trained RoBERTa-Large for NLI task and includes the combination of SNLI Bowman et al. (2015), MNLI Williams et al. (2018), FEVER Nie et al. (2019) and ANLI Nie et al. (2020) datasets as the training data. Hence, this incurs a potential data leakage problem, as we also test it on FEVER and ANLI. To some extent, it explains the strong performance of RoBERTa on FEVER and ANLI in Table LABEL:tab:human_results.
The model is used as a classifier in evaluation, and three classes are available. Class 0 corresponds to “entail”, class 1 corresponds to “neutral” and class 2 means “not entail”. In our experiment setting, we only regard class 0 as “Yes” label and combine the remaining two classes to be “No” label.
Entailer
Tafjord et al. (2022) Entailer is a T5-based model Raffel et al. (2020) trained to answer hypotheses by building proof trees containing chains of reasoning. It can either generate valid premises for a given hypothesis or predict a score for a given premise and hypothesis. We evaluate Entailer-11B (model name “allenai/entailer-11b” on HuggingFace) in our experiments. Similarly, Entailer dataset Dalvi et al. (2021) is in the training set, making this model very competitive when evaluating on the same dataset.
We strictly follow the official implementation of the model to acquire class labels. [^4] https://github.com/allenai/entailment_bank/blob/main/entailer.md The “entailment_verifier” is called to decide if the hypothesis can be implied from the premise. If the answer is “True”, then the class label will be “Yes”. And vice versa.
[…]
[sections/tables.tex]promptSensitivity
[…]
[sections/tables.tex]fewshotResults
B.2 Instruction-Finetuned models
These are the more recent “general-purpose” language models trained on a collection of NLP tasks described using instructions, leading to generalization abilities to solve unseen tasks described using new instructions. Following are models included in our evaluation.
Flan-T5-xxl
Chung et al. (2022) It is instruction-tuned from T5 Raffel et al. (2020) on 1.8K+ tasks. We adopt a publicly available version on HuggingFace with model card name “google/flan-t5-xxl”. Flan-T5-xxl is also exposed to data leakage issues. BoolQ Clark et al. (2019), ECQA Aggarwal et al. (2021) and ANLI Nie et al. (2020) have appeared in its training data. However, this is not a serious problem in the finetuning stage because we transform original datasets into entailment verification format before using them for model finetuning.
We extract labels from the model by focusing on the output probabilities of two words, “Yes” and “No”. After applying the softmax function to those two probabilities, we finalize the label as the word with a probability larger than a given threshold.
GPT-3.5
Brown et al. (2020) It is a general-purpose autoregressive decoder-only LMs accessible via the OpenAI Completions API. [^5] https://platform.openai.com/docs/api-reference/completions We utilize “text-davinci-003” in OpenAI’s API for evaluation, and the label determination procedure is quite similar to the one in Flan-T5-xxl. We monitor the probability of “Yes” and “No” tokens. After applying the softmax function to them, we select the token that attains a probability larger than a threshold to be our label.
GPT-4
OpenAI (2023) It is the latest generative model published by OpenAI which is optimized for creativity and long context inputs. It is accessible via the OpenAI Chat API. [^6] https://platform.openai.com/docs/api-reference/chat We adopt plain “gpt-4” in OpenAI’s API for our experiments. Unlike other models, the output probabilities are not accessible. Alternatively, we firstly constrain the model to predict “Yes” or “No” only by putting some instructions into the prompt. Then, we check the output text. In the ideal case, we can directly use the output text as the label since only “Yes” or “No” is produced. Otherwise, we need to randomly choose one as the label.
B.2.1 Prompts for Robustness Assessment
To assess the robustness of the model in section 2.2, we design different prompt formats but hold the order of premise and hypothesis in the prompt unchanged. Table LABEL:tab:prompt_sensitivity presents all prompts we tested with Flan-T5-xxl and their corresponding averaged results across datasets. The results suggest that the model is insensitive to the variation in the prompt and yields relatively consistent results. This characteristic is maintained in Flan-T5-xxl + Class and Flan-T5-xxl + Rank as well since they are generally constructed based on Flan-T5-xxl.
B.2.2 Few-Shot Evaluation
Few-shot is an effective and promising strategy when testing the performance of a model Brown et al. (2020). We also include this analysis by randomly picking two examples from Entailer as demonstrations and incorporating them into the prompt. We test Flan-T5-xxl, Flan-T5-xxl + Rank, and GPT-4 in this setting and represent results in Table LABEL:tab:fewshot_results. The few-shot setting yields promising improvement for Flan-T5-xxl, substantiating that the few-shot is a beneficial approach to teaching the prompt to the model. However, it is not as helpful as our finetuning strategies, which give even better performance. On the other hand, few-shot does not bring significant gains for Flan-T5-xxl + Rank, suggesting that finetuning has already helped the model have a comprehensive understanding of the prompt, and extra demonstrations are unnecessary. As for the GPT-4, simply using examples from Entailer and applying the same prompt for all datasets seem detrimental.
[…]
[sections/tables.tex]majorityResults
Appendix C Majority Prediction and Label Imbalance
In Table LABEL:tab:majority_results, we show the performance of an oracle model that predicts the most frequent label in a dataset. For a label-balanced dataset, the macro-F1 score of such a majority prediction model would be 0.67 (precision 0.5 and recall 1.0). The datasets in the evaluation set have some label imbalance, as evidenced by the lower majority label prediction scores. Since we convert existing 3-class NLI and multi-choice QA datasets into our binary classification task format, it inherently has more not support labels. We have more support instances for the rationale datasets since the dataset creators usually only annotate the rationale for the right choice. Specifically, the ECQA dataset only has positive instances, leading to a 1.0 macro-F1 score for majority prediction (support label). Since it has all support labels, any model predicting even a single non support label gets penalized severely, as is seen in ECQA results. Because of this label imbalance in the datasets, we report the macro-F1 scores instead of accuracy or micro-F1.
Appendix D Results on Hypothesis Groundedness
We define the groundedness of a hypothesis based on the type of knowledge required to establish its validity. Using this categorization, we classify the datasets into two categories as follows:
  • Knowledge-based: A knowledge-based hypothesis typically contains entity-grounded or commonsense knowledge. Such a hypothesis is usually well-defined (i.e., the premise is not required to ground it), and models can potentially leverage their internal knowledge to verify the entailment without using the premise. The datasets that contain such knowledge-based hypotheses are FEVER, ANLI, BoolQ, Entailer, and ECQA.
  • Contextual: A contextual hypothesis requires localized knowledge, i.e., the premise/context grounds the hypothesis. These hypotheses cannot be verified on their own without using the premise. Datasets with contextual hypotheses are WaNLI, CosQA, SIQA, DREAM, and RACE.
[…]
[sections/figures.tex]groundedness
Performance w.r.t. groundedness
Figure LABEL:fig:hypothesis_groundedness shows the aggregate performance across the two hypothesis categories described above. First, we observe that performance on the knowledge-based hypotheses is higher on average than contextual hypotheses for all models (except GPT-3.5 and GPT-4, for which they are comparable). This is likely because models can potentially leverage their internal knowledge stored in the parameters to check the validity of the knowledge-based hypotheses. Such a scenario is harder for contextual cases since the hypothesis solely depends on the context and would require a strong context understanding to predict the hypothesis entailment.
The models in the plot are arranged with increasing complexity, i.e., along the x-axis, either the model size or the amount/variety of pretraining data increases. We observe that the performance on contextual hypotheses almost consistently improves with scaling, whereas the trend is not so strong for the knowledge-based case. This shows that scaling is useful for training LLMs in understanding contexts. However, this does not necessarily improve LLMs’ knowledge retention or retrieval skills.
[…]
[sections/figures.tex]reasoningAnalysis1
Appendix E Human Alignment Experiments
We adopted Amazon Mechanical Turk (MTurk) Crowston (2012) for data collection. Two annotation formats (Figure LABEL:fig:eval_annotate and Figure LABEL:fig:reason_annotate) were devised for the human evaluation task and reasoning type analysis task, respectively. During the annotations, each annotator was compensated according to \$15/hour per the U.S. minimum wage.
E.1 Human Evaluation Details
In each HIT, the annotator was presented with a format exactly like Figure LABEL:fig:eval_annotate, including detailed task descriptions and label explanations. Annotators were expected to read the premise and claim first, then determine the supportiveness of the claim based on the premise and choose the corresponding label. Initially, we only provided three labels — “support”, “irrelevant” and “contradict”. But later, we realized that annotators could not explicitly identify labels for some ambiguous instances where the premise only partially supported or contradicted the claim. Hence, we introduced two weak labels (“partially support” and “partially contradict”) to remedy this issue. When collating results for analysis, we internally combined “support” and “partially support” to be “support”, and the rest to be “not support”, aligning to the standard entailment verification (EV) setup. Each instance was annotated by 3 MTurk annotators, and a majority verdict determined the label. The Fleiss’s kappa score Fleiss (1971) we got was 0.6, indicating a moderate level of agreement among annotators.
[…]
[sections/figures.tex]evalAnnotate
E.2 Reasoning Type Annotation Details
In every HIT, we used Figure LABEL:fig:reason_annotate as the reasoning type annotation format. It was an intense job with two tasks involved. The instructions and label explanations were explicitly stated at the beginning of the format. For task 1, after reading the premise and the claim, annotators should decide whether the supportiveness of the claim could be decided by only referring to the information in the premise. If yes, annotators needed to choose the corresponding difficulty level of reasoning about the supportiveness of the claim in task 2. Otherwise, the type of missing information should be decided in task 1, and task 2 did not apply to these cases. As stated in 2.5, we aggregated answers from two tasks and categorized them into four types. Both R1 and R2 were types where the premise contained all necessary information. The difference was that the difficulty level of reasoning for R1 was Easy while for R2 was Moderate. We combined Missing Entity-grounded Information and Missing Commonsense Information to be R3. Finally, the Missing Localized Information label corresponded to R4.
Given that some NLP background knowledge was required to understand task descriptions and label explanations, we recruited Computer Science graduate students instead of the general public to finish this MTurk annotation. Similarly, every instance was assigned to 2 students, and the majority vote determined the label. The Fleiss’s kappa score for this job was 0.62, showing a substantial inter-annotator agreement.
[…]
[sections/figures.tex]reasonAnnotate
E.3 Reasoning Type Examples
Table LABEL:tab:reasoning_examples incorporates three examples from each reasoning type, providing more insight into those types.
In the third example of R1, the first sentence in the premise states that “iron oxide” comes from “oxygen” and “rust”. The second sentence shows those two substances are “gases” at room temperature. Therefore, combining them will be sufficient to entail the hypothesis.
However, unlike R1, the third example of R2 requires three steps of reasoning that “De Baandert was a multi-use stadium.”, “It was mostly used for football matches.” and “The stadium was able to hold 22,000 people.”. The hypothesis can be disproved with those steps because “22,000 people” is just the maximum capacity.
As for the first example of R3, some missing commonsense information like “it is not wise to give more money to a person who keeps playing in a detrimental situation.” should be combined with the premise to disprove the hypothesis.
In the first example of R4, the next movement of “Addison” is the missing information specific to the context depicted by the premise. The hypothesis can not be directly disproved without that piece of information.
E.4 Reasoning Type Analysis
We depict the aggregated results of the reasoning type annotation for each dataset in Figure LABEL:fig:reasoning_analysis1. Here, the first row shows the frequency of each reasoning type in that dataset, and the corresponding plot in the second row compares the human and GPT-4 macro-F1 scores. We fade out the columns with data percentages less than 5% because such low frequency might not lead to conclusive observations.
We note that type R1 is most prominent in Entailer dataset. Here, we observe that humans are significantly better than models. This shows that humans are usually more consistent with simple deductive reasoning. Similar findings about consistency in human deductive reasoning skills have been reported in prior works Sanyal et al. (2022); Nguyen et al. (2023).
Reasoning type R2, which requires more complex reasoning, is dominant in ANLI and CosQA. For this type, we find that models are superior to humans. Complex reasoning requires two skills: understanding multiple relevant information in the premise and then using them for reasoning. We hypothesize that models are stronger context processors than humans because they have been trained on long-context data OpenAI (2023).
The reasoning type R3 is present in ANLI, CosQA, and SIQA. From Table LABEL:tab:dataset_comparisons, we know that ANLI mostly require entity-grounded knowledge, whereas CosQA and SIQA specifically test commonsense knowledge. Here, we find that humans are stronger than models in commonsense knowledge (CosQA and SIQA), whereas models are better in ANLI that requires entity-grounded knowledge. This shows that humans can infer missing social/commonsense knowledge more easily since these are inherently known to humans. In contrast, models can retrieve the entity-grounded knowledge stored in their parameters more efficiently.
Lastly, we find that the reasoning type R4, indicating missing localized knowledge, is also prominent in ANLI, CosQA, and SIQA. Here, we find that the trends are a bit mixed. We find that for SIQA, humans are better at recognizing missing localized knowledge, but in CosQA, models outperform humans. This is likely because SIQA typically contains short contexts based on everyday social situations that are easier for humans to understand. In contrast, CosQA has longer contexts with rarer situations requiring more complex understanding.
Overall, we conclude that GPT-4 outperforms humans in complex deductive reasoning and situations involving entity-grounded knowledge, whereas humans are more consistent at simple reasoning and situations requiring commonsense knowledge.
Appendix F Finetuning LLMs
In this section, we describe more details about the training dataset used for finetuning, our negative data collection strategy that was used in the ranking formulation, and other finetuning details.
F.1 Training Dataset Selection
To train the Flan-T5-xxl model, we create a training dataset using representative datasets from each category. We pick the ANLI, RACE, and ECQA datasets to represent NLI, contextual QA, and rationale categories, respectively. We note that the amount of training data is quite low for the Rationale category. We also include the StrategyQA Geva et al. (2021) dataset in the training set to alleviate this. Similar to BoolQ, StrategyQA has a Yes/No type of questions and their corresponding explanations. We convert the question-answer pair into a hypothesis using the QA-to-statement converter Chen et al. (2021) as described in Section 2.1.
For a given premise and a valid hypothesis, generate five alternate hypotheses contradicted by the premise. Try to avoid using the negation words such as “not”, “never”, etc. The output should be numbered from 1 to 5. Premise: {premise} Hypothesis: {hypothesis}
List of Boxes 2 Prompt format for generating alternate negative hypothesis for a given premise-hypothesis pair. Please refer to Section F.2 for details.
[…]
[sections/tables.tex]cotResults
F.2 Negative Data Collection for Ranking
In the ranking formulation, for a given premise and hypothesis pair $(p,h)$, we need to find some weaker hypothesis $h^{\prime}$ to use the ranking loss defined by Equation 2. We collect such weaker hypotheses in two ways and then combine them to form the training data. The two techniques are described below:
  • Using incorrect options: The contextual QA category has naturally occurring negative data. For a given question and choices, we pair the hypothesis corresponding to the correct option with all other hypotheses corresponding to the wrong options to create the ranked data.
  • GPT-3.5 prompting: The other way we generate negative data is by prompting GPT-3.5. Specifically, we use the prompt format shown in Box 2 to generate alternate hypotheses contradicted by the original premise. We only select premise and hypothesis pairs that originally have support label. GPT-3.5 generated hypotheses are then considered negative samples and paired with the original hypothesis. We repeat this for all the training datasets (ANLI, RACE, ECQA, and StrategyQA).
F.3 Hyperparameters and other details
During training, we select the learning rate from the set $\{7e^{-5},1e^{-4},2e^{-4}\}$, per GPU batch size from the set ${6,8}$, margin $m$ in Equation 2 from the set $\{0.2,0.3,0.5\}$, and warmup ratio $0.1$. The model is trained for 1400 steps on a cluster of 8 A6000 GPUs. We evaluate the model every 200 steps and save the checkpoint if the model shows improvements on a held-out development set.
Appendix G Chain-of-Thought Filtering
We study three variants of CoT Filtering as mentioned below:
  • $\mathcal{B}+\textrm{SC}$: This is the self-consistency baseline. Here, $\mathcal{B}$ is the base model used to sample CoTs. We sample 40 CoTs for each instance before computing the majority predicted label.
[…]
Following Wang et al. (2023), we use four different base CoT model: UL2 Tay et al. (2023), Codex-001 Brown et al. (2020), LaMDA-137B Thoppilan et al. (2022), and ChatGPT OpenAI (2022). Further, we compute the CoTs and analyze the performance of the above methods for three multi-choice QA datasets, namely, CommonsenseQA Talmor et al. (2019) and AI2 Reasoning Challenge Clark et al. (2018) (easy (ARC-e) and challenge (ARC-c) variants). Please refer to Wang et al. (2023); Wei et al. (2022) and the associated code [^7] https://openreview.net/attachment?id=1PL1NIMMrw&name=supplementary_material for details on the CoT prompt formats. The results are shown in Table LABEL:tab:cot_results. We note a consistent improvement between the three variants, with Flan-T5-xxl + Rank model performing the best. This demonstrates the advantage of our entailment finetuning approach. Please refer to Section 4 for more findings.
G.1 Filtering Examples
Table LABEL:tab:filter_examples presents three CoT reasoning examples, each including two outputs that are kept and three that are filtered out by ranking. According to the table, outputs supported by strong rationales are ranked highly and kept. On the other hand, if the rationale is irrelevant to the prediction (like rationale 3 in example 1), the rationale itself is incomplete (like rationale 5 in example 1), or the rationale supports another option rather than the prediction (like rationale 4 in the example 1), then such output will have a low entailment score leading to a lower ranking and getting filtered out.
[…]
[sections/tables.tex]reasoningExamples
[…]
[sections/tables.tex]filterExamples

### v3

@@
-Minds versus Machines: Rethinking Entailment Verification with Language Models
+Are Machines Better at Complex Reasoning? Unveiling Human-Machine Inference Gaps in Entailment Verification
 Abstract
-Humans make numerous inferences in text comprehension to understand discourse.
-This paper aims to understand the commonalities and disparities in the inference judgments between humans and state-of-the-art Large Language Models (LLMs).
-Leveraging a comprehensively curated entailment verification benchmark, we evaluate both human and LLM performance across various reasoning categories.
-Our benchmark includes datasets from three categories (NLI, contextual QA, and rationales) that include multi-sentence premises and different knowledge types, thereby evaluating the inference capabilities in complex reasoning instances.
-Notably, our findings reveal LLMs’ superiority in multi-hop reasoning across extended contexts, while humans excel in tasks necessitating simple deductive reasoning.
-Leveraging these insights, we introduce a fine-tuned Flan-T5 model [^1] https://huggingface.co/soumyasanyal/entailment-verifier-xxl that outperforms GPT-3.5 and rivals with GPT-4, offering a robust open-source solution for entailment verification.
-As a practical application, we showcase the efficacy of our finetuned model in enhancing self-consistency in model-generated explanations, resulting in a 6% performance boost on average across three multiple-choice question-answering datasets.
-Minds versus Machines: Rethinking Entailment Verification with Language Models
+Making inferences in text comprehension to understand the meaning is essential in language processing.
+This work studies the entailment verification (EV) problem of multi-sentence premises that requires a system to make multiple inferences implicitly.
+Studying EV for such complex premises is important because modern NLP problems, such as detecting inconsistent model-generated rationales, require complex multi-hop reasoning.
+However, current textual inference datasets mostly contain short premises that only partially focus on these challenges.
+To address this, we compile an EV benchmark that includes datasets from three NLP domains (NLI, contextual QA, and rationales) containing multi-sentence premises.
+On benchmarking humans and LLMs, we find that LLMs are better than humans in multi-hop reasoning across extended contexts, while humans perform better in simple deductive reasoning tasks.
+We also finetune a Flan-T5 model [^1] https://huggingface.co/soumyasanyal/entailment-verifier-xxl for EV using two training objectives to obtain a strong open-source model that outperforms GPT-3.5 and rivals GPT-4.
+Finally, we use this model to filter out inconsistent model-generated rationales in self-consistency decoding, resulting in a 6% accuracy improvement on average across three MCQ datasets.
+Are Machines Better at Complex Reasoning? Unveiling Human-Machine Inference Gaps in Entailment Verification
 Soumya Sanyal1111 Tianyi Xiao1111 Jiacheng Liu2222 Wenya Wang3333Xiang Ren1111 $1$University of Southern California $2$University of Washington $3$Nanyang Technological University, Singapore soumyasa@usc.edu
@@
 A prevailing notion in cognitive psychology exists that humans make numerous inferences to understand discourse and text Garnham (1989).
-These inferences, serving diverse functions in text comprehension, extend beyond resolving lexical ambiguities or determining the referent of a pronoun; they play a crucial role in linking information from disparate sections of a text to establish its literal meaning.
-Cognitive studies Buschman et al.
-(2011); Cowan (2001) have shown that an average human brain has a limited capacity to retain only four chunks in short-term memory, indicating a limitation of human inference abilities.
-Similarly, prior works Gururangan et al.
-(2018); McCoy et al.
-(2019); Geiger et al.
-(2020); Clark et al.
-(2020); Sanyal et al.
-(2022) have shown that large language models (LLMs) are still subpar at understanding negations in inference, use spurious correlations while reasoning, cannot reason through multi-step compositions robustly, etc.
-With the advent of large language models (LLMs) and their applications on natural language inference (NLI) Minaee et al.
-(2021); Nie et al.
-(2020), understanding the key difference in the inference abilities between humans and LLMs is crucial to making further progress, which is currently missing.
-In this work, we aim to address this by first evaluating both humans and current LLMs Liu et al.
-(2019); Tafjord et al.
+These inferences play a crucial role in linking information from disparate sections of a text to establish its literal meaning and are closely associated with reasoning.
+With the recent applications of Large Language Models (LLMs) Devlin et al.
+(2019); Raffel et al.
+(2020); Brown et al.
 (2020); Chung et al.
-(2022); Brown et al.
-(2020); OpenAI (2023) on a comprehensive entailment verification benchmark and then analyzing the performance across different reasoning categories based on the ease of inference deduction and the type of knowledge required.
-Existing textual inference datasets such as SNLI Bowman et al.
+(2022) in NLP tasks that require inference skills Minaee et al.
+(2021), it is thus essential to understand and improve upon the limitations of LLMs concerning different aspects of language inferences.
+In this work, we focus on the task of entailment verification (EV) that classifies whether a given context supports a hypothesis.
+To assert the validity of a hypothesis, a system has to make multiple inferences from the given context and its internal knowledge, which requires complex multi-hop reasoning.
+[^2] In a multi-hop reasoning instance, a system has to infer implicit inferences by combining information from the premise, to predict the support of a hypothesis.
+Verifying the entailment of such complex premises has important applications in rationale-generating LLMs, such as chain-of-thoughts (CoT) Wei et al.
+(2022), that typically generate multi-sentence rationales with incomplete information and suffer from inconsistent reasoning Ye and Durrett (2022).
+While EV is similar to Natural Language Inference (NLI) Dagan et al.
+(2006); Manning and MacCartney (2009), some interesting challenges distinguish it from NLI.
+Many existing textual inference datasets, such as SNLI Bowman et al.
 (2015), MNLI Williams et al.
-(2018), etc., mostly contain short sentence premises that do not completely encapsulate the challenges of multi-sentence, complex reasoning.
-Predicting the entailment of such complex premise-hypothesis pairs often requires multi-hop reasoning, inferring missing information, robustness to spurious correlation, etc., which are largely missing from simpler datasets Gururangan et al.
+(2018), etc., mostly contain short sentence premises that only partially encapsulate the challenges of multi-sentence premises requiring complex reasoning.
+Predicting the entailment of such complex premise-hypothesis pairs often requires multi-hop reasoning, inferring missing information, etc., which is lacking from standard NLI datasets Gururangan et al.
 (2018); McCoy et al.
 (2019).
-Also, real-world scenarios that require inference skills, such as understanding stories, dialogues, etc., usually contain multi-sentence premises.
-Therefore, we select multiple datasets across three categories (NLI, contextual QA, and rationales) and convert them into NLI format, as required, to create the evaluation benchmark.
-As shown in Table LABEL:tab:dataset_comparisons, the datasets used in our study typically contain multi-sentence premises that require different types of knowledge to predict the entailment.
-Our analyses show that LLMs are stronger than humans at entailment tasks that involve multi-hop reasoning across long contexts.
-This suggests that training LLMs to reason on diverse long context data is beneficial for entailment verification on complex cases.
-In contrast, we find that humans are better at solving entailment tasks that require simple deductive reasoning using substitutions, negations, etc., indicating that current LLMs still lack consistency along these reasoning aspects.
-Further, we find that humans and LLMs perform comparably in entailment tasks requiring inferring some missing knowledge.
-These findings are further depicted in Figure LABEL:fig:motivation with motivating examples.
-Among the different LLMs, we find that general instruction-finetuned models are better than task-finetuned models trained on a specific dataset category.
-We leverage this finding by finetuning a Flan-T5 Chung et al.
-(2022) model on a training subset containing datasets from each category.
-We explore two different training approaches: a classification-based finetuning that learns to directly predict the label and a ranking-based finetuning that learns to rank the most supported hypothesis from a given pair of hypotheses for a given premise.
-We find that ranking-based finetuning is superior to classification as it can learn a softer decision boundary.
-Overall, our fine-tuned models outperform GPT-3.5 and perform comparably to GPT-4 on the benchmark, thus providing a strong open-sourced model for entailment verification.
+Thus, studying EV in the context of modern LLMs is essential and is currently missing.
+To this end, we first compile an evaluation benchmark to study the EV problem by selecting multiple datasets across three categories: NLI, contextual QA, and rationales.
+As shown in Table LABEL:tab:dataset_comparisons, the datasets typically contain multi-sentence premises that require inferring different types of knowledge to predict the entailment.
+Thus, this benchmark is more complex than standard NLI datasets and can be used as a new evaluation benchmark.
+Next, we evaluate LLMs and humans on this benchmark and make some interesting observations.
+Cognitive studies Buschman et al.
+(2011); Cowan (2001) have shown that an average human brain has a limited capacity to retain only four chunks in short-term memory, indicating a limitation of human inference abilities over long contexts.
+Our analysis shows that LLMs are indeed stronger than humans at tasks that involve multi-hop reasoning across long contexts, reinforcing this human limitation.
+In contrast, humans are better in cases that require simple deductive reasoning using substitutions, negations, etc., indicating that current LLMs lack consistency along these reasoning aspects.
+Further, humans and LLMs perform comparably in instances requiring inferring missing knowledge.
+These findings are depicted in Figure LABEL:fig:motivation with motivating examples.
+Additionally, on comparison between LLMs, we find that models finetuned on a specific dataset category are usually strong within the category but don’t generalize well on unseen categories.
+In contrast, instruction-finetuned models are better on average, with the best performing model, GPT-4, achieving 0.79 macro-F1 and outperforming the best-open-sourced model Flan-T5-xxl by 0.08 macro-F1 on average.
+In order to bridge this gap, we finetune a Flan-T5 Chung et al.
+(2022) model on a training subset containing datasets from each category of the above data collection.
+To this end, we explore two different training approaches: a classification-based finetuning that learns to directly predict the label and a ranking-based finetuning that learns to rank the most supported hypothesis from a given pair of hypotheses for a given premise.
+Ranking-based finetuning is better than classification, specifically in contextual QA datasets, as it can learn a softer decision boundary.
+Overall, our fine-tuned Flan-T5 outperforms GPT-3.5, baseline Flan-T5, and performs comparably to GPT-4 on the benchmark, thus providing a strong open-sourced model for entailment verification.
 Finally, we demonstrate the utility of our finetuned models on a downstream application of filtering unfaithful model-generated explanations.
-Recent works Wei et al.
-(2022); Zhou et al.
-(2023); Wang et al.
-(2023) have prompted LLMs to generate natural language explanations for model predictions, but such reasoning steps can sometimes be inconsistent, i.e., the explanation does not entail the model’s final prediction Ye and Durrett (2022).
-In self-consistency (SC) Wang et al.
-(2023), multiple reasoning paths are first sampled from the LLM decoder for a given instance and then aggregated to predict the most consistent answer.
-We use our finetuned models to filter out non-entailed reasoning chains in the SC decoding strategy before aggregating the final prediction, which leads to 6% performance improvement on average across three MCQ datasets.
-2 Humans vs. LLMs on Entailment Verification
-In this section, we define the task of entailment verification (EV), datasets used to create the evaluation benchmark, the evaluation procedure for evaluating different LLM baselines and humans, and compare them across different reasoning types.
-[…]
+Self-consistency (SC) Wang et al.
+(2023) decoding samples multiple model-generated reasoning paths from the LLM decoder and aggregates them to predict the most consistent answer.
+We use our finetuned Flan-T5 to filter out non-entailed reasoning chains before aggregating the final prediction, which leads to 6% performance improvement on average across three MCQ datasets.
+Overall, our contributions can be summarized as follows:
+  1. 1.
+    Using existing resources, we develop a benchmark for Entailment Verification (EV) of multi-sentence premises for both training and evaluation purposes.
+  2. 2.
+    We compare and contrast humans and GPT-4 on this benchmark and conclude that humans are more robust on simple deductive reasoning while LLMs excel at complex inferences.
+  3. 3.
+    We use a ranking-based objective to finetune Flan-T5-xxl model for the EV task, achieving comparable performance to GPT-4 and demonstrate the utility of our model in filtering non-entailed CoT rationales.
+2 Entailment Verification using LLMs
+In this section, we formally define the task of entailment verification (EV), the datasets used to create the evaluation benchmark, and the evaluation procedure for evaluating different LLM baselines.
+For a given premise $p$ and a hypothesis (or claim) $h$, the task of entailment verification is to determine whether the context has information that directly confirms the hypothesis or not, i.e., whether the hypothesis follows from the information present in the context.
+This is a binary classification task defined as $f(p,h)=\{\textrm{{support}},\textrm{{not support}}\}$, where $f$ is a classifier (human/LLM).
 2.1 Evaluation Benchmark
-To benchmark humans and LLMs on the task of entailment verification, a basic criterion for dataset selection is that the dataset should contain a premise, a hypothesis, and a support/not support label associated with it.
-Additionally, we list some more desirable characteristics we want to cover in this study, as discussed below:
+In this section, we list some desirable properties we want to include in the EV benchmark, followed by details of the dataset categories we select.
   • Type of Premise: Typically, NLI datasets, such as SNLI, MNLI, etc., do not contain more than one sentence in the premise, potentially leading to shortcut learning.
@@
 This information is often missing from the premise and has to be inferred implicitly.
-Lastly, localized information is all other knowledge provided for understanding the events, people, or items mentioned in the premise that is not grounded to any known entity.
+Lastly, localized information is all other knowledge provided for understanding the events, people, or items mentioned in the premise that are not grounded to any known entity.
 This information depends on the premise’s specific context and, thus, is impossible to infer unless stated explicitly.
@@
 Please refer to Appendix A for more details on the datasets used.
-[…]
-[sections/tables.tex]humanResults
 Natural Language Inference
-NLI is an obvious choice of data source.
-While converting an NLI dataset for our task, we merge the neutral and contradict labels to the not support label.
+Given the close connection between NLI and EV, it is an obvious choice to consider appropriate NLI datasets for the benchmark.
+To convert an NLI dataset for EV, we merge the neutral and contradict labels to the not support label.
 We use the following NLI datasets in our benchmark: WaNLI Liu et al.
@@
 (2020).
+[…]
+[sections/tables.tex]humanResults
 Contextual QA
-Next, we consider QA datasets where the task is to answer a question based on a given context and some options.
-We use an off-the-shelf QA-to-statement converter model Chen et al.
+Next, we consider multiple-choice question-answering datasets where the task is to answer a question based on a given context and some options.
+We use a QA-to-statement converter model Chen et al.
 (2021) to generate a hypothesis statement for each question option pair.
-Then, the hypothesis corresponding to the correct choice is marked as “support”, while the rest are marked as “not support” to create the entailment verification dataset.
-Overall, we include the following datasets from this category: Cosmos QA (CosQA) Huang et al.
+Then, the hypothesis corresponding to the correct choice is marked as “support”, while the rest are marked as “not support”.
+This process is depicted in Figure LABEL:fig:contextual_qa in Appendix A.
+We include the following datasets from this category: Cosmos QA (CosQA) Huang et al.
 (2019), SocialIQA (SIQA) Sap et al.
@@
 Rationale
-Lastly, we consider data sources where human-annotated explanations are available that justify the original hypothesis (or the correct option, in the case of QA datasets).
+Lastly, we consider data sources where human-annotated explanations are available that justify the original hypothesis.
 In this case, we use the rationales as the premise.
@@
 (2021).
+2.2 Evaluation Metric
+We use the macro-F1 score as the primary evaluation metric for comparing LLMs on the entailment verification task because there are label imbalances in our evaluation datasets.
+The macro-F1 score computes the unweighted mean of F1 scores for each class, ensuring equal importance for each class irrespective of the label statistics.
+Please refer to Appendix C for more discussions on the label imbalance of each dataset.
 2.3 LLM Evaluation Setup
-We evaluate two types of LLMs on the task of entailment verification, as categorized below:
+We evaluate two types of LLMs on the task of EV, as categorized below:
 Task-finetuned LLMs
-In this category, the models considered are already finetuned for either NLI or the exact entailment verification task itself.
-We evaluate RoBERTa (finetuned on NLI data) and Entailer-11B (finetuned on Entailer data) in this category.
-Please refer to Appendix B.1 for more details on the evaluation setup for these models.
+In this, the models considered are already finetuned on some subset of the benchmark.
+We evaluate two models: RoBERTa Liu et al.
+(2019) (finetuned on NLI datasets) and Entailer-11B Tafjord et al.
+(2022) (finetuned on Entailer dataset).
+Refer to Appendix B.1 for more details on the evaluation setup for these models.
+Premise: {premise} Hypothesis: {hypothesis} Question: Given the premise, is the hypothesis correct? Answer:
+List of Boxes 1 Prompt used to evaluate instruction-finetuned LLMs for entailment verification.
 Instruction-finetuned LLMs
-These are “general-purpose” language models trained on a collection of NLP tasks described using instructions, leading to generalization abilities to solve unseen tasks described using new instructions.
-Here, we evaluate Flan-T5-xxl, GPT-3.5, and GPT-4 models.
+These are language models trained on a collection of NLP tasks described using instructions, leading to generalization abilities to solve unseen tasks described using new instructions.
+Here, we evaluate Flan-T5-xxl Chung et al.
+(2022), GPT-3.5 Brown et al.
+(2020), and GPT-4 OpenAI (2023) models.
 To compute the label, we first modify a given premise-hypothesis pair $(p,h)$ into a prompted input $\mathcal{P}$ using the prompt template as shown in Box 1.
@@
 For GPT-4 evaluation, we directly check for the “Yes” / “No” label prediction as the token probabilities are not accessible via the API.
-Please refer to Appendix B.2 for more details about the models and ablations on few-shot prompts.
-Premise: {premise} Hypothesis: {hypothesis} Question: Given the premise, is the hypothesis correct? Answer:
-List of Boxes 1 Prompt used to evaluate instruction-finetuned LLMs for entailment verification.
-2.3 Evaluation Metric
-We use the macro-F1 score as the primary evaluation metric for comparing LLMs on the entailment verification task because there are label imbalances in our evaluation datasets.
-The macro-F1 score computes the unweighted mean of F1 scores for each class, ensuring equal importance for each class irrespective of the label statistics.
-Please refer to Appendix C for more discussions on the label imbalance of each dataset.
-2.4 Human vs. LLMs
-First, we randomly sample 100 instances for each dataset (i.e., 1000 instances in total) and conduct a human evaluation on this subset to estimate average human performance.
+Please refer to Appendix B.2 for more details about the models and ablations on prompts.
+3 Evaluation of Humans and LLMs
+First, we randomly sample 100 instances from each dataset (i.e., 1000 instances in total) and conduct a human evaluation on this subset to estimate average human performance.
 Please refer to Appendix E.1 for more details on the annotation procedure.
@@
 Table LABEL:tab:human_results shows the overall evaluation results.
-Among LLMs, we observe that instruction-finetuned LLMs are stronger than task-finetuned models that are trained on some selected datasets.
-This shows that using a specific dataset for fine-tuning does not necessarily improve the general inference abilities of LLMs.
-On comparisons between humans and LLMs, we find that humans beat all the baseline LLMs, except GPT-4.
-This shows that existing open-sourced LLMs are not at par with humans on this task.
-Additionally, although humans and GPT-4 perform comparably on average, there are large misalignments in different individual datasets.
+3.1 Comparison among LLMs
+We observe that the task-finetuned models (rows 1-2 in Table LABEL:tab:human_results) are weaker on average compared to the instruction-finetuned models (rows 3-5 in Table LABEL:tab:human_results).
+However, there are some interesting exceptions.
+RoBERTa, which is finetuned on NLI datasets, performs at par with GPT-4 on FEVER and outperforms it on WaNLI but falls behind on contextual QA and rationale datasets.
+On the other hand, Entailer-11B, which is finetuned on the Entailer dataset, outperforms GPT-4 on Entailer but lags on most of the other datasets.
+This demonstrates that finetuning an LLM using datasets from one of these categories is ineffective in outperforming models trained on more general data.
+Overall, we observe that GPT-4 is the best-performing entailment verification model on average and Flan-T5-xxl is the best open-source LLM for this task.
+3.2 Comparison between Humans and LLMs
+In Table LABEL:tab:human_results, we report the human performance and the corresponding annotation agreement [^3] We use a pairwise agreement ratio that computes the fraction of matched annotations over all pairs of annotations.
+(in parenthesis) for each dataset.
+We note that the agreement ratios are lowest for Contextual QA datasets, especially CosQA and SIQA.
+Questions in these datasets are often based on commonsense scenarios, and sometimes, the support for the wrong hypothesis can be debatable.
+Please refer to Appendix D for examples.
+Between humans and LLMs, we find that humans beat all the baseline LLMs, except GPT-4.
+This shows that existing open-sourced LLMs are subpar with humans on this task.
+Additionally, although humans and GPT-4 perform comparably on average, we note that large misalignments exist in different individual datasets.
 Specifically, we observe that ANLI, CosQA, SIQA, and Entailer are the four datasets with $>0.1$ absolute macro-F1 difference.
-Next, we perform more analysis on these four datasets to better understand these misalignments.
+We analyze these four datasets in Section 3.3.
 […]
 [sections/figures.tex]reasoningAnalysis2
-2.5 Reasoning Type Analysis
-We design a new analysis to further understand the reasons behind such misalignment between models and humans.
-First, we categorize the type of reasoning required to successfully predict an entailment into the following four categories:
-  • Simple Deductive (R1): In this type, the premise contains sentences that can be minimally combined in one step to predict the support for the hypothesis.
-This type typically tests skills such as substitution, negation, synonyms, etc.
-  • Complex Deductive (R2): In this, more than one step of reasoning using information from the premise is required to solve the task.
-Typically, this type tests skills like mathematical reasoning, using multiple information in context to arrive at a conclusion, etc.
-  • Missing Entity-grounded/Commonsense Knowledge (R3): This is a case where some essential commonsense or entity-grounded knowledge is missing in the premise.
-This gap can potentially be bridged by the model’s parametric knowledge obtained from pretraining.
-  • Missing Localized Knowledge (R4): In this, information very specific to the context of the premise is missing.
-Typically, this is information about the subjects in the context and is not grounded to any entities that can be known via the internet.
-It is practically impossible for the model (or even humans) to infer such missing information.
-Please refer to Appendix E.3 for examples of each reasoning type.
-Given this categorization, we first annotate the reasoning type of the 100 previously sampled instances for each dataset with absolute macro-F1 difference $>0.1$.
-We make this choice intending to attribute the largely misaligned datasets (namely, ANLI, CosQA, SIQA, and Entailer) since some random noise in the annotation and sampling process can potentially also cause some misalignment.
-Please refer to Appendix E.2 for more details on the annotation setup.
-Figure LABEL:fig:reasoning_analysis2 depicts the aggregated results for each reasoning type.
-The top plot shows the frequency of each reasoning type, and the bottom plot compares the human and GPT-4 macro-F1 scores.
-Please refer to Appendix E.4 for a detailed analysis of individual datasets.
+[…]
+[sections/tables.tex]finetuneResults
+3.3 Effect of Reasoning Type
+We design an analysis to understand further the misalignments between GPT-4 and humans. First, we categorize the type of reasoning required to predict an entailment into the following four categories:
+  • Simple Deductive (R1): The premise contains sentences that can be minimally combined in one step to predict the support for the hypothesis.
+This typically tests skills such as substitution, understanding negations, word meanings, etc.
+  • Complex Deductive (R2): More than one step of reasoning is required to solve the task. Typically, this tests skills like mathematical reasoning, combining multiple information in context, etc.
+  • Missing Entity-grounded/Commonsense Knowledge (R3): In this, some essential commonsense or entity-grounded knowledge is missing in the premise.
+Both humans and LLMs can implicitly invoke such information from the memory or the parametric knowledge obtained from pertaining, respectively.
+  • Missing Localized Knowledge (R4): In this, information very specific to the premise is missing.
+Typically, this is information about the subjects in the context and is not grounded in any known entities.
+It is practically impossible for humans or LLMs to infer such information.
+We note that these categories are mutually exclusive.
+[^4] Deductive reasoning implies that the premise has all the necessary information.
+Thus, any missing knowledge instance falls under inductive reasoning.
+Figure LABEL:fig:reasoning_analysis2 depicts the aggregated results.
+The top plot shows the percentage of each reasoning type among 400 samples from ANLI, CosQA, SIQA, and Entailer, and the bottom plot compares the human and GPT-4 macro-F1 scores.
+Please refer to Appendices E.2, E.3, and E.4 for details on the annotation setup, examples of each reasoning type, and detailed analysis, respectively.
 The first type in Figure LABEL:fig:reasoning_analysis2 is simple deductive reasoning ($\sim$ 25% data).
 Here, humans perform better than GPT-4 by a small margin.
-Instances that require simple deductive reasoning usually use substitutions, negations, paraphrasing, etc., to prove entailment (refer to Table LABEL:tab:reasoning_examples for examples).
-We find that humans are more robust than GPT-4 in performing such simple deductive reasoning tasks, as observed in prior works Sanyal et al.
+Instances that require simple deductive reasoning usually use substitutions, negations, paraphrasing, etc., to prove entailment.
+We find that humans are more robust than GPT-4 in performing such simple deductive reasoning tasks, which is also observed in prior works Sanyal et al.
 (2022); Nguyen et al.
 (2023).
-[…]
-[sections/tables.tex]finetuneResults
-Next, we find that GPT-4 significantly outperforms the human baseline on complex reasoning that constitutes $\sim$ 20% data.
-This type of reasoning usually requires two skills: understanding multiple relevant information in the premise and combining them for reasoning.
-There are two potential reasons for this misalignment.
-First, compared to humans, GPT-4 is likely a stronger context processor, especially for long premises, since it has been trained on long-context data sources OpenAI (2023).
-Alternatively, another potential confounding factor is that AMT workers are typically incentivized to quickly finish annotations to maximize earnings.
-Thus, instances that require multi-step reasoning can have noisier annotations, potentially underestimating true human performance.
+Next, we find that GPT-4 significantly outperforms the humans on complex reasoning that constitutes $\sim$ 20% data.
+This usually requires two skills: understanding multiple relevant information in the premise and combining them for reasoning.
+GPT-4 is likely a stronger context processor, especially for long premises, since it has been trained on long-context data sources OpenAI (2023).
+In contrast, cognitive studies Buschman et al.
+(2011); Cowan (2001) have shown that an average human brain can retain only four chunks in short-term memory, thus limiting the long-context processing abilities of humans.
 Lastly, we observe that approximately 30% of the data has some missing entity-grounded or commonsense information while $\sim$ 25% of the data has missing localized information.
-To correctly predict entailment in such instances, a system should be able to infer some of the missing grounded knowledge while not hallucinating specific localized information not mentioned in the premise.
+To correctly predict entailment in such cases, a system needs to infer some missing grounded knowledge while not hallucinating specific localized information not mentioned in the premise.
 We find that both humans and models are comparable across the reasoning types R3 and R4, with R4 being more challenging.
@@
 4 Training LLMs for Entailment Verification
-In Section 2.4, we observed that open-sourced LLMs are lacking in performance compared to humans and close-sourced models such as GPT-4.
-Thus, in this section, we propose training strategies to finetune an open-sourced model that can perform competitively w.r.t.
-GPT-4.
-To this end, we finetune Flan-T5-xxl model using the training data from some of the above-mentioned datasets.
-Specifically, we choose the training portions of ANLI, RACE, and ECQA for finetuning the model.
+In Section 3.2, we observed that open-sourced LLMs lack performance compared to humans and close-sourced models such as GPT-4.
+From Section 3.1, we know that task-finetuned models perform well on a category when finetuned on data from the same category.
+Using this insight, we finetune the Flan-T5-xxl model on the train splits of datasets from each category, resulting in using ANLI Nie et al.
+(2020), RACE Lai et al.
+(2017), and ECQA Aggarwal et al.
+(2021) for finetuning.
 Please refer to Appendix F.1 for more details on our training dataset selection criteria.
-We describe two approaches to finetuning and then discuss our findings.
-[…]
-[sections/figures.tex]cotFlowchart
+Next, we describe the finetuning approaches and our key findings.
 4.1 Finetuning Formulations
-This section describes the two fine-tuning formulations explored in this work.
+This section describes the two finetuning formulations explored in this work.
 Classification
@@
 We refer to this finetuned model as “Flan-T5-xxl + Class”.
+[…]
+[sections/figures.tex]cotFlowchart
 Ranking
 In this approach, for a given premise and hypothesis pair $(p,h)$, we define a weaker hypothesis $h^{\prime}$ as a statement such that the premise $p$ supports $h$ more strongly than $h^{\prime}$.
-Then, for a given triple $(p,h,h^{\prime})$, we formulate the ranking task as predicting the hypothesis that is more supported by the premise.
+Then, for a given triplet $(p,h,h^{\prime})$, we formulate the ranking task as predicting the hypothesis that is more supported by the premise.
 Given the triplet $(p,h,h^{\prime})$, we define the margin ranking loss as follows:
@@
 This can lead to better generalization, especially for contextual QA datasets.
-Sometimes, the wrong choice can be relatively less favorable w.r.t.
-to the best option in QA instead of being absolutely incorrect.
+Sometimes, the wrong choice can be relatively less favorable compared to the best option in QA instead of being incorrect.
+As discussed in Section 3.2, this is indicated by the low agreement between human raters.
 Training to hard-classify the hypothesis for such options can be avoided by ranking them with the best hypothesis (corresponding to the right choice), thus learning a softer classification boundary.
@@
 For our models, we separately average the results for the datasets already seen in training (namely, ANLI, RACE, and ECQA) and unseen during training into two columns, seen and unseen, respectively.
-First, we observe that all our finetuned models are consistently better across nine out of ten datasets than the baseline Flan-T5-xxl.
-Finetuning improves 0.07 macro-F1 on average.
+First, we observe that all our finetuned models are consistently better across nine of ten datasets than the baseline Flan-T5-xxl.
+Finetuning improves 0.07 macro-F1 on average over Flan-T5-xxl.
 This shows that finetuning is overall beneficial in training the model on the task of entailment verification.
@@
 Thus, Flan-T5-xxl + Rank is a strong open-sourced model for entailment verification and can be used as an alternative to GPT-4.
+[…]
+[sections/figures.tex]cot
 5 Application: Filtering CoT Rationales
@@
 (2022) in LLMs, whereby multiple CoT rationales are sampled for a given input instance and a majority voting overall predicted labels is considered as the final prediction.
-But generative LLMs can potentially output rationales that are inconsistent Ye and Durrett (2022), i.e., the rationale does not support the corresponding model prediction.
+However, generative LLMs can potentially output rationales that are inconsistent Ye and Durrett (2022), i.e., the rationale does not support the corresponding model prediction.
 Such inconsistency can, in turn, degrade the overall self-consistency results.
-This motivates an important application of entailment verification systems in this paradigm.
+Evaluating the consistency between an explanation and the corresponding prediction can be framed as an entailment verification (EV) task, as described below.
 Please refer to Appendix G.2 for examples of consistent and inconsistent CoTs.
 Approach
-As shown in Figure LABEL:fig:cot_application, we can use a verifier as an intermediate step before computing the majority vote, whereby the inconsistent rationales are filtered out.
-We achieve this by first scoring all the rationales (premises) and their corresponding predictions (hypothesis [^2] For MCQ datasets, typically, the prediction is one of the choices.
-We use Chen et al.
-(2021) as defined in Section 2.1 to convert the question-answer pair into a statement.) using Equation 1.
-Then, we use the top-$k$ rationales for majority voting.
-We set $k=10$ for all our experiments.
-[…]
-[sections/figures.tex]cot
+As shown in Figure LABEL:fig:cot_application, we can use a verifier as an intermediate filtering step to filter out the inconsistent rationales before computing the majority vote.
+For this, we define the generated CoT rationale as the premise and use the QA-to-statement model Chen et al.
+(2021) as defined in Section 2.1 to convert the question and model’s prediction into a hypothesis.
+Next, we calculate the entailment score of all the premise-hypothesis pairs using a verifier (Equation 1).
+Finally, we select the top-$k$ rationales for majority voting, discarding the rest.
+We set $k=5$ for all our experiments.
 Findings
@@
 (2020), LaMDA-137B Thoppilan et al.
-(2022), and ChatGPT OpenAI (2022)).
+(2022), and ChatGPT [^5] ChatGPT refers to the gpt-3.5-turbo-0613 model.
+OpenAI (2022)).
 Please refer to Appendix G for details on the datasets and more comparisons with Flan-T5-xxl.
@@
 For instance, filtering UL2 generated-rationales can even achieve comparable performance with vanilla SC over LaMDA-137B.
-In comparison, the gains for filtering ChatGPT CoTs are $\sim$ 1.5 %.
+In comparison, the gains for filtering ChatGPT CoTs are $\sim$ 1.7 %.
 This shows that weaker models are prone to generating inconsistent CoTs and thus benefit more from this approach.
@@
 Reasoning in LLMs
-With the advent of strong, general-purpose LLMs Brown et al.
+With the advent of general-purpose LLMs Brown et al.
 (2020); Chung et al.
@@
 While the progress is impressive, some of these are either specialized approaches for math-specific datasets or heavily rely on close-sourced LLMs (GPT-3.5, GPT-4, etc.) for verification.
-In contrast, here we focus on natural language datasets and develop a strong open-sourced LLM that can be easily deployed to check entailment errors in LLM reasoning chains.
+In contrast, here we focus on natural language datasets and develop an open-sourced model for the verification task.
+Ranking Objective
+Prior works have explored the benefits of ranking objective both from a theoretical perspective Narasimhan and Agarwal (2013) and specifically for NLP classification tasks Li et al.
+(2019); Briakou and Carpuat (2020).
+Our ranking-based objective is inspired by these works that find that margin-based loss can be beneficial in training plausibility estimation models.
 7 Conclusion
-We studied the entailment verification problem in the context of LLMs.
-Specifically, we sourced datasets across three different categories (NLI, contextual QA, and rationales) and analyzed both human and model performance across these datasets.
-We found some misalignments between the two, whereby models are better than humans in complex reasoning and humans are relatively more consistent on simpler reasoning tasks.
-We also explored different finetuning objectives to train LLMs for verification.
-Our fine-tuned models outperform GPT-3.5 and are at par with GPT-4.
-Finally, we demonstrated a practical application of entailment verification in filtering out inconsistent model-generated rationales while using self-consistency decoding.
-Overall, we believe our work provides a meaningful comparison of inference capabilities between humans and LLMs, a comprehensive evaluation benchmark, an open-sourced entailment verification model, and interesting applications of the task.
+We studied the EV problem in the context of LLMs.
+Specifically, we sourced datasets across three different categories (NLI, contextual QA, and rationales) and analyzed human and LLM performance on the benchmark.
+We found that LLMs are better than humans in complex reasoning, while humans are more consistent on simpler reasoning tasks.
+We also explored different training objectives to finetune open-sourced LLMs for EV.
+Our finetuned model outperforms GPT-3.5, baseline Flan-T5-xxl, and is comparable to GPT-4.
+Finally, we applied the EV model to filter out inconsistent model-generated CoTs in self-consistency decoding, achieving improvements over the baseline self-consistency approach.
 Limitations
 Even though our work demonstrates exciting results on entailment verification tasks by finetuning LLMs, several limitations can be potentially improved.
+Data Processing
+Our strategy to convert a QA pair into a statement using the QA-to-statement converter model can have errors that can cascade both in the evaluation dataset and our fine-tuned models.
+Better strategies for this pipeline would help with data quality.
+Finetuning
 We only tried encoder-decoder-based models for finetuning.
-However, other models with different architectures (like decoder-only) can also be considered.
-Our strategy to convert a QA pair into a statement using the QA-to-statement converter model can have errors that can cascade both in the evaluation dataset and our fine-tuned models.
+However, other models with different architectures (like decoder-only) can also be considered in the future.
 For computing the entailment score in Equation 1, we only considered the probability of “Yes” and “No” tokens in the entire vocabulary.
@@
 Answer generation as a training objective can be more robust since it is a stricter objective than our scoring technique.
+Potential Risks
+We using existing datasets, with some post-processing, to train our models.
+Thus, any issues in the existing dataset in terms of bias, toxicity, etc., can potentially affect our model.
+Also, since we use a pretrained checkpoint, we also inherit any existing biases in the baseline model.
+The model is trained to always output scores, even if the data is well outside the training distribution.
+This is an existing issue with most NLP models and can be mitigated by additional checks for domain shift.
 Acknowledgments
-This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the DARPA MCS program under Contract No.
-N660011924033, the Defense Advanced Research Projects Agency with award W911NF-19-20271, NSF IIS 2048211, and gift awards from Google and Amazon.
+This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the Defense Advanced Research Projects Agency with award HR00112220046, and NSF IIS 2048211.
 The views and conclusions contained herein are those of the authors.
@@
 While converting an NLI dataset for our task, we merge the neutral and contradict labels to the not support label.
+[^6] We note that merging the labels for the NLI datasets in this benchmark is valid, since the underlying meaning of the ‘neutral’ label is ‘no sufficient evidence to support’ in all three datasets.
+NLI datasets that don’t follow this label definitions cannot be modified in this manner.
 We use the following NLI datasets in our benchmark:
@@
 (2018).
-And the test split is used when doing evaluation.
+The test split is used when doing the evaluation.
   • FEVER Nie et al.
@@
 Premises in this dataset typically contain multiple sentences, which is one of our focus areas.
-As the test split is not available, we report results on dev split for evaluation.
+As the test split is unavailable, we report results on the dev split for evaluation.
   • ANLI Nie et al.
@@
 During evaluation, the test split is considered.
+[…]
+[sections/figures.tex]contextualQA
 Contextual QA datasets
@@
 Then, the hypothesis corresponding to the correct choice is marked as “support”, while the rest are marked as “not support” to create the entailment verification dataset.
+We depict this process in Figure LABEL:fig:contextual_qa.
+The green box is the valid hypothesis and the red ones are the invalid hypothesis corresponding to the given context.
 Overall, we include the following datasets from this category:
@@
 The key challenge in this dataset is understanding people’s everyday narratives described in the context that can have some missing commonsense knowledge that needs to be inferred implicitly.
-Since the test split is missing, we evaluate models on dev split instead.
+Since the test split is missing, we evaluate models on the dev split instead.
   • SocialIQA (SIQA) Sap et al.
@@
 This dataset has more nuanced commonsense knowledge requirements, which makes it a challenging dataset for our task.
-Similarly, results on dev split are reported given test is missing.
+Similarly, results on the dev split are reported, given that the test is missing.
   • DREAM Sun et al.
@@
 Questions in this dataset require complex and difficult entailment-like inference to solve, making it a good set for evaluation.
-The test split is lacking for this dataset and we can only report results on dev split.
+The test split is lacking for this dataset, and we can only report results on the dev split.
   • RACE Lai et al.
@@
 (2021) and crowdsourcing.
-The dataset mainly contains science domain statements and tests simple deductive reasoning skills whereby sentences from the premise have to be combined to either support or refute the hypothesis.
-The test split is also missing for this dataset and we can only evaluate on the dev split.
+The dataset mainly contains science domain statements and tests simple deductive reasoning skills whereby sentences from the premise must be combined to either support or refute the hypothesis.
+The test split is also missing for this dataset, and we can only evaluate it on the dev split.
   • ECQA Aggarwal et al.
@@
 We only use the explanations for the correct choice as the explanations for the incorrect choices are often trivial.
-It is a complete dataset and, by convention, we use the test split for evaluation.
+It is a complete dataset and by convention, we use the test split for evaluation.
 Appendix B Details on LLM Evaluation
@@
 (2020); Liu et al.
-(2019) This is a strong pre-trained RoBERTa-Large model with corresponding model card on HuggingFace [^3] https://huggingface.co/models called “ynie/roberta-large-snli_mnli_fever_anli_R1_R2_R3-nli”.
+(2019) This is a strong pre-trained RoBERTa-Large model with the corresponding model card on HuggingFace [^7] https://huggingface.co/models Wolf et al.
+(2020b) called “ynie/roberta-large-snli_mnli_fever_anli_R1_R2_R3-nli”.
 It is a specifically pre-trained RoBERTa-Large for NLI task and includes the combination of SNLI Bowman et al.
@@
 Hence, this incurs a potential data leakage problem, as we also test it on FEVER and ANLI.
-To some extent, it explains the strong performance of RoBERTa on FEVER and ANLI in Table LABEL:tab:human_results.
+To some extent, it explains the strong performance of RoBERTa on NLI datasets in Table LABEL:tab:human_results.
 The model is used as a classifier in evaluation, and three classes are available.
 Class 0 corresponds to “entail”, class 1 corresponds to “neutral” and class 2 means “not entail”.
-In our experiment setting, we only regard class 0 as “Yes” label and combine the remaining two classes to be “No” label.
+In our experiment setting, we only regard class 0 as the “Yes” label and combine the remaining two classes to be the “No” label.
 Entailer
@@
 Similarly, Entailer dataset Dalvi et al.
-(2021) is in the training set, making this model very competitive when evaluating on the same dataset.
+(2021) is in the training set, making this model very competitive when evaluating the same dataset.
 We strictly follow the official implementation of the model to acquire class labels.
 [^8] https://github.com/allenai/entailment_bank/blob/main/entailer.md The “entailment_verifier” is called to decide if the hypothesis can be implied from the premise.
-If the answer is “True”, then the class label will be “Yes”.
-And vice versa.
+If the answer is “True”, then the class label will be “Yes” and vice versa.
 […]
@@
 These are the more recent “general-purpose” language models trained on a collection of NLP tasks described using instructions, leading to generalization abilities to solve unseen tasks described using new instructions.
-Following are models included in our evaluation.
+The models included in our evaluation from this category are described below.
 Flan-T5-xxl
@@
 (2020) on 1.8K+ tasks.
-We adopt a publicly available version on HuggingFace with model card name “google/flan-t5-xxl”.
+We adopt a publicly available version on HuggingFace Wolf et al.
+(2020a) with model card name “google/flan-t5-xxl”.
 Flan-T5-xxl is also exposed to data leakage issues.
@@
 (2019), ECQA Aggarwal et al.
-(2021) and ANLI Nie et al.
+(2021), and ANLI Nie et al.
 (2020) have appeared in its training data.
@@
 We extract labels from the model by focusing on the output probabilities of two words, “Yes” and “No”.
-After applying the softmax function to those two probabilities, we finalize the label as the word with a probability larger than a given threshold.
+After applying the softmax function to those two probabilities, we finalize the label as the word with a probability larger than a given threshold, as described in Equation 1.
 GPT-3.5
 Brown et al.
-(2020) It is a general-purpose autoregressive decoder-only LMs accessible via the OpenAI Completions API.
-[^5] https://platform.openai.com/docs/api-reference/completions We utilize “text-davinci-003” in OpenAI’s API for evaluation, and the label determination procedure is quite similar to the one in Flan-T5-xxl.
-We monitor the probability of “Yes” and “No” tokens.
-After applying the softmax function to them, we select the token that attains a probability larger than a threshold to be our label.
+(2020) It is a general-purpose autoregressive decoder-only LLM accessible via the OpenAI Completions API.
+[^9] https://platform.openai.com/docs/api-reference/completions We utilize “text-davinci-003” in OpenAI’s API for evaluation, and the label determination procedure is similar to the one in Flan-T5-xxl.
 GPT-4
-OpenAI (2023) It is the latest generative model published by OpenAI which is optimized for creativity and long context inputs.
+OpenAI (2023) It is the latest generative model published by OpenAI, which is optimized for creativity and long context inputs.
 It is accessible via the OpenAI Chat API.
@@
 Unlike other models, the output probabilities are not accessible.
-Alternatively, we firstly constrain the model to predict “Yes” or “No” only by putting some instructions into the prompt.
-Then, we check the output text.
-In the ideal case, we can directly use the output text as the label since only “Yes” or “No” is produced.
-Otherwise, we need to randomly choose one as the label.
-B.2.1 Prompts for Robustness Assessment
+Thus, we constrain the model to predict a single token as its generated prediction.
+In the ideal case, we can directly use the output text as the label if “Yes” or “No” is produced.
+If some other token is generated, we choose one of the labels at random.
+We note that this occurrence is rare in our GPT-4 evaluation runs.
+B.2.1 Prompt Robustness Evaluation
 To assess the robustness of the model in section 2.3, we design different prompt formats but hold the order of premise and hypothesis in the prompt unchanged.
 Table LABEL:tab:prompt_sensitivity presents all prompts we tested with Flan-T5-xxl and their corresponding averaged results across datasets.
-The results suggest that the model is insensitive to the variation in the prompt and yields relatively consistent results.
-This characteristic is maintained in Flan-T5-xxl + Class and Flan-T5-xxl + Rank as well since they are generally constructed based on Flan-T5-xxl.
+The results suggest that the Flan-T5-xxl model is robust to the variation in the prompt and yields relatively consistent results.
+This characteristic is maintained in Flan-T5-xxl + Class and Flan-T5-xxl + Rank as well since they are developed over based on Flan-T5-xxl.
 B.2.2 Few-Shot Evaluation
@@
 Since it has all support labels, any model predicting even a single non support label gets penalized severely, as is seen in ECQA results.
-Because of this label imbalance in the datasets, we report the macro-F1 scores instead of accuracy or micro-F1.
-Appendix D Results on Hypothesis Groundedness
-We define the groundedness of a hypothesis based on the type of knowledge required to establish its validity. Using this categorization, we classify the datasets into two categories as follows:
-  • Knowledge-based: A knowledge-based hypothesis typically contains entity-grounded or commonsense knowledge.
-Such a hypothesis is usually well-defined (i.e., the premise is not required to ground it), and models can potentially leverage their internal knowledge to verify the entailment without using the premise.
-The datasets that contain such knowledge-based hypotheses are FEVER, ANLI, BoolQ, Entailer, and ECQA.
-  • Contextual: A contextual hypothesis requires localized knowledge, i.e., the premise/context grounds the hypothesis.
-These hypotheses cannot be verified on their own without using the premise.
-Datasets with contextual hypotheses are WaNLI, CosQA, SIQA, DREAM, and RACE.
+Because of this label imbalance in the datasets, we report the macro-F1 scores instead of accuracy or micro-F1 metrics.
+Appendix D Debatable Cases in Contextual QA
+We include two examples of SIQA dataset in Table LABEL:tab:contextual_examples to illustrate the debatable hypotheses in Contextual QA datasets.
+These examples demonstrate that some less probable hypotheses can also be considered plausible in certain situations.
+In the first example, option “run screaming” best describes the audience’s reactions after listening to a scary ghost story.
+However, there are also situations where people are so terrified and start to cry, making the option “cry” somehow debatable.
+Similarly, in the second example, it is common sense that people will express gratefulness after someone else builds a house for them.
+But it is reasonable to say that Quinn is intrigued by the house after seeing the amazing architecture.
+In both cases, apart from the most adequate one, there are still options leading to some partially supported hypotheses, which explains the low human annotation agreement in these cases.
 […]
-[sections/figures.tex]groundedness
-Performance w.r.t. groundedness
-Figure LABEL:fig:hypothesis_groundedness shows the aggregate performance across the two hypothesis categories described above.
-First, we observe that performance on the knowledge-based hypotheses is higher on average than contextual hypotheses for all models (except GPT-3.5 and GPT-4, for which they are comparable).
-This is likely because models can potentially leverage their internal knowledge stored in the parameters to check the validity of the knowledge-based hypotheses.
-Such a scenario is harder for contextual cases since the hypothesis solely depends on the context and would require a strong context understanding to predict the hypothesis entailment.
-The models in the plot are arranged with increasing complexity, i.e., along the x-axis, either the model size or the amount/variety of pretraining data increases.
-We observe that the performance on contextual hypotheses almost consistently improves with scaling, whereas the trend is not so strong for the knowledge-based case.
-This shows that scaling is useful for training LLMs in understanding contexts.
-However, this does not necessarily improve LLMs’ knowledge retention or retrieval skills.
+[sections/tables.tex]contextualExamples
 […]
 [sections/figures.tex]reasoningAnalysis1
-Appendix E Human Alignment Experiments
+Appendix E Human Evaluation and Analysis
 We adopted Amazon Mechanical Turk (MTurk) Crowston (2012) for data collection.
@@
 Each instance was annotated by 3 MTurk annotators, and a majority verdict determined the label.
-The Fleiss’s kappa score Fleiss (1971) we got was 0.6, indicating a moderate level of agreement among annotators.
+To ensure annotation quality, we conducted two rounds of qualification for annotators using the finalized template.
+In the first round, we used ten questions from the datasets and evaluated 400 mTurk annotators.
+We retained 100 annotators from this batch with an accuracy greater than 70%.
+Among remained ones, we repeated this process for another ten questions and selected annotators with an accuracy greater than 80%.
+Finally, we retained 35 annotators who had annotated the human evaluation set.
+And the Fleiss’s kappa score Fleiss (1971) we got was 0.6, indicating a moderate level of agreement among annotators.
 […]
@@
 E.2 Reasoning Type Annotation Details
+We annotate the reasoning type of the 100 sampled instances for each dataset with absolute macro-F1 difference $>0.1$ in Table LABEL:tab:human_results.
+We make this choice intending to attribute the largely misaligned datasets (namely, ANLI, CosQA, SIQA, and Entailer) since some random noise in the annotation and sampling process can potentially also cause some misalignment.
 In every HIT, we used Figure LABEL:fig:reason_annotate as the reasoning type annotation format.
-It was an intense job with two tasks involved.
-The instructions and label explanations were explicitly stated at the beginning of the format.
-For task 1, after reading the premise and the claim, annotators should decide whether the supportiveness of the claim could be decided by only referring to the information in the premise.
+The instructions and label explanations were explicitly stated at the beginning of the HIT.
+For task 1, after reading the premise and the claim, annotators had to decide whether the supportiveness of the claim could be decided by only referring to the information in the premise.
 If yes, annotators needed to choose the corresponding difficulty level of reasoning about the supportiveness of the claim in task 2.
-Otherwise, the type of missing information should be decided in task 1, and task 2 did not apply to these cases.
-As stated in 2.5, we aggregated answers from two tasks and categorized them into four types.
+Otherwise, the type of missing information had to be decided in task 1, and task 2 did not apply to these cases.
+We aggregated answers from the two tasks and categorized them into four types.
 Both R1 and R2 were types where the premise contained all necessary information.
@@
 Finally, the Missing Localized Information label corresponded to R4.
-Given that some NLP background knowledge was required to understand task descriptions and label explanations, we recruited Computer Science graduate students instead of the general public to finish this MTurk annotation.
-Similarly, every instance was assigned to 2 students, and the majority vote determined the label.
+Given that some strong NLP background knowledge was required to understand task descriptions and label explanations, we recruited Computer Science graduate students instead of the general mTurk workers to finish this annotation.
+Every instance was assigned to 2 students, and the majority vote determined the label.
 The Fleiss’s kappa score for this job was 0.62, showing a substantial inter-annotator agreement.
@@
 Therefore, combining them will be sufficient to entail the hypothesis.
-However, unlike R1, the third example of R2 requires three steps of reasoning that “De Baandert was a multi-use stadium.”, “It was mostly used for football matches.” and “The stadium was able to hold 22,000 people.”.
+However, unlike R1, the third example of R2 requires three steps of reasoning that “De Baandert was a multi-use stadium.”, “It was mostly used for football matches.”, and “The stadium was able to hold 22,000 people.”.
 The hypothesis can be disproved with those steps because “22,000 people” is just the maximum capacity.
@@
 We pick the ANLI, RACE, and ECQA datasets to represent NLI, contextual QA, and rationale categories, respectively.
+We select the datasets with diverse entailment challenges and aim to maximize the total training data.
 We note that the amount of training data is quite low for the rationale category.
-We also include the StrategyQA Geva et al.
+Thus, we also include the StrategyQA Geva et al.
 (2021) dataset in the training set to alleviate this.
@@
 F.3 Hyperparameters and other details
-During training, we select the learning rate from the set $\{7e^{-5},1e^{-4},2e^{-4}\}$, per GPU batch size from the set ${6,8}$, margin $m$ in Equation 2 from the set $\{0.2,0.3,0.5\}$, and warmup ratio $0.1$.
+During training, we select the learning rate from the set $\{7e^{-5},1e^{-4},2e^{-4}\}$, per GPU batch size from the set $\{6,8\}$, margin $m$ in Equation 2 from the set $\{0.2,0.3,0.5\}$, and warmup ratio $0.1$.
 The model is trained for 1400 steps on a cluster of 8 A6000 GPUs.
@@
 […]
-[…]
+  • +Flan-T5-xxl + Rank + $\mathcal{B}+\textrm{{Flan-T5-xxl + {Rank}}{} + SC}$: This is the same as above, but instead, we use our ranking-finetuned Flan-T5-xxl model for filtering.
 Following Wang et al.
@@
 Please refer to Section 5 for more findings.
-G.1 Filtering Examples
+[…]
+[sections/figures.tex]k-selection
+G.1 Ablation of Top-k Filtering
+We examine our top-$k$ selecting strategy by choosing different values of $k$ and compare their results.
+We pick up $k$ from $\{3,5,10,20,30\}$ and present the corresponding aggregated results in Figure LABEL:fig:k-selection.
+The general trend is that, as $k$ increases, the model performance increases first and then starts to decrease, reaching the peak at $k=5$.
+When $k$ is small, there are only a limited number of CoT rationals, and the majority voting process is vulnerable to potential outliers.
+On the other hand, if k is too large, then many noisy results are included, leading to poorer performance.
+G.2 Examples of Filtered CoTs
 Table LABEL:tab:filter_examples presents three CoT reasoning examples, each including two outputs that are kept and three that are filtered out by ranking.
 According to the table, outputs supported by strong rationales are ranked highly and kept.
-On the other hand, if the rationale is irrelevant to the prediction (like rationale 3 in example 1), the rationale itself is incomplete (like rationale 5 in example 1), or the rationale supports another option rather than the prediction (like rationale 4 in the example 1), then such output will have a low entailment score leading to a lower ranking and getting filtered out.
+On the other hand, if the rationale is irrelevant to the prediction (like rationale 3 in example 1), the rationale itself is incomplete (like rationale 5 in example 1), or the rationale supports another option rather than the prediction (like rationale 4 in the example 1), then such output has a low entailment score leading to a lower ranking and getting filtered out.
 […]